New ZOOM room:

News:

L1 accepted by MB - Kaori PPD - Donato: TSG - Piparo: computing confirmed - Canelli: Phys Coord https://indico.cern.ch/event/1025576/contributions/4306380/attachments/2224312/3767105/MB236_SP.pdf CMS: ppl worried because we are loosing part of the float due to the HGCAL. (fortunately) Atlas is behind, so we will need check with them and CERN, clearly we are expecting delays. Funding is quite critical: India, Turkey, Russia seem no to be able or willing to pay (order of 10M). Work on LS2 seems to be fine. Expecting to be ready for TB end of the year. Safety training: has to be done by all of us online.

Framework & production

No news. No machine added to the cloud since March. By the end of the month all nodes will be in condor.

Analysis

B physics

Bs->tautau: meeting las week. Ppl interested in this analysis, students from India interested. Hevjin has only 1 additional year. Enrico: finishing the work on the second tagger, need to do that because of the new trigger. Alberto's activity is well defined, will restart asap. Lepton Flavor Universality: CMS has 2 analyses, rD* and rK.

Heavy composite neutrinos

Relaunched the processing. Final application of the systs from Muon POG. Taken new ele objects, this should make some difference in the low mass (Not in the signal region).

VBS

Improvements in the analysis, Matteo not working directly on that anymore. Possible opening for a dim-6 operator interpretation. Work for students.

ML

INFN organizing a Hackaton in June. @ CERN there was a meeting among ppl interested in ML to have CERN joining Ellis.

Mu regression

working on the features to use.

WH_SS

No news.

H/Z->JPsi gamma

Rocco: signal generation progressing. Swiched to ultra legacy version, rerun the analysis, seems slightly worse. optimization re

Higgs

no news

EPR

Detector

Trigger

no news

Timing Layer

2 activities going on. cooling of the plate: system has bee studied, needs update cooling with TEC: jigs ready for tests on the TECs.

DT

Anna: short presentation yesterday about how to use Fluka simulation of the background rate. The Fluka info are particle fluxes, not easy to translate these info into rates for each detector. Ppl responsible for the background simulation checking. Gif: the situation is not clear. We need a test beam on the chamber before any further irradiation, in July. Some new wires in the chamber showed NO sign of ageing.

Tracker/tracking

A.O.B.

-- RobertoRossin - 2021-04-13

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2021-04-13 - RobertoRossin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback